

********************************************************************
UARS WEEKLY STATUS REPORT - 18 JUN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this
past week without apparent problems.
We are evaluating the recently reprocessed V20 products before their
release to the DAAC.
***********
PEM
***********
PEM/AXIS was active from June 2 through June 6 and half of the instrument
was operated from June 13 through June 17. During this time, PEM/AXIS
generated deposition profiles from electron precipitation which were inverted
from measured X-rays.
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
HALOE was powered on June 15, 1999. All engineering parameter are nominal
and science channel outputs are within expected operating range.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF would like to say "BEST WISHES" to Al Marshall of the HRDI group
in his retirement.
The UK Meteorological Office has resumed sending UKMO data to the CDHF.
Day gaps have been received and data is currently up to date.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days CLA (7.0) PROD (L1-L3) (461R-48C=413L)
0 days GPI (4.05) PROD (MLS V5 3AL) (39R-0C-1F=39L)
58 days WIN (5.11) PROD (L1-L3) (1540R-735C-4F=801L)
CDHF will load April 1999 current day DAAC data in June 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 June 1999 (Orbit 42249 , GMT Day 156) through 11 June
(GMT Day 162).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 157/0916 42269 AXIS Off
HALOE 158/1205 42286 Off
WINDII 160/2330 42323 Off
A Forward-to-Reverse Yaw Around maneuver (number 79) was successfully
executed in Orbit 42348 beginning 162/14:00:34 and completing at
120/14:37:44 on 11 June 1999. The SSPP Instruments (ACRIM, SOLSTICE,
SUSIM) were enabled as well as HRDI and PEM in day only operations. Due to
the SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 50 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 23.8 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 24.0 and 24.5 percent respectively.
The next (reverse-to-forward) YAW is scheduled for 10 July 1999 (DOY 191).
The beta angle ranged between the angles of 27.5 degrees to 0.8 degrees for
this report period. The beta angle is now increasing toward a maximum
angle of 34 degrees on 26 June 1999 (DOY 177).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.7 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 04 June 1999 and 11 June 1999 are provided in Appendix
D.
The clock error ranged from -7.1 to -3.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 09
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% WINDII 01
% HALOE 01
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/130 - 2287.496913
Transponder B: 99/151 - 2287.499087
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 13 minutes, 05.396 seconds (an increase of 2.176 seconds since
the last report period). This is a 0.0358 percent data loss which equals a
99.9642 percent data capture for the mission.
The increase of 2.176 seconds occurred over one (1) day:
7 2.176 seconds on Day 157 due to no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed until a new FORMATS software
release can be made to accommodate the special Y2K planning aids.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The line fill test has been postponed indefinitely since all instrument
PIUs stated their instruments can take data at lower altitudes.
MEETINGS
FOT attended the monthly status meeting hosted by Ed Macie on Wednesday 09
June 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (reverse-to-forward) is scheduled for 10
July 1999 (DOY 191).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 June 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42269 157/0916 99-175 PEM Power OFF, AXIS 1 & 2
42286 158/1205 99-199 HALOE Power OFF
42323 160/2330 99-200 WINDII Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 June 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIQBVOLTAGE Y-LO WINDII Quiet Bus Voltage 42255-6
HRPSEQBIMON Y-HI HRDI PSE Quiet Bus Current Mon 42255-6,61-2
WIPBVOLTAGE Y-LO WINDII Pulse Bus Voltage 42276-7
WIQBVOLTAGE Y-LO WINDII Quiet Bus Voltage 42276
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 42276-7
OBSVLBEON Y-LO Load Bus Voltage EON 42286
OBSBATVEON3 Y-LO Battery 3 voltage at EON 42286
WIQBVOLTAGE Y-LO WINDII Quiet Bus Voltage 42320
HRPSEQBIMON Y-HI HRDI Quiet Bus Current Mon 42320
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 June 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-No clock adjusts for this report period
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 June 1999
04 June 1999 - GMT Day 155 Beta = 27.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-112.0 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-33.6 +6.76/+5.15 +31.2/-15.2 24.0 1.022 22.5
3 +33.6/-33.6 +5.47/+4.19 +30.4/-16.4 23.8 1.022 22.3
11 June 1999 - GMT Day 162 Beta = 0.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -95.2/-117.6 -3.05/-4.94 +0.0/ +0.0 16.5 0.000 0.0
2 +67.2/-5.6 +5.47/+4.19 +30.4/-15.2 23.8 1.015 21.6
3 +28.0/-39.2 +4.51/+2.92 +29.6/-16.0 23.8 1.015 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 June 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT 42255 156/1106 14170 Generic Late Acquisition
99-028 FOT 42308 159/2311 14170 Generic Late Acquisition
99-029 FOT 42317 160/1405 20954 Late Acq. PTP reset
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT see list June 99 14170 Generic Late Acq
99-029 FOT 42317 160/1405 20954 Late Acq. PTP reset
SEVNTFW99/24